A number is 6 greater than 1/2 of an of the number. If the sum of the numbers is 21, find the numbers.
Answer by fractalier(6550)

Call the numbers x and y. Then we have
x = 6 + (1/2)y and
x + y = 21
Substitute the first into the second and get
6 + (1/2)y + y = 21
6 + (3/2)y = 21
(3/2)y = 15
Now multiply by 2/3 and get
y = 10
x is then 11.